OpenStudy (anonymous):

PLEASE HELP! IN ANY WAY YOU CAN!! Andrew makes $6 an hour plus $9 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime. (3 points) Part B: Create an equation that shows the amount of wages earned, S, for working y hours of overtime. Hint: Remember to include in the equation the amount earned from working 40 hours. (3 points) Part C: Andrew earned $330 in 1 week. How many hours (regular plus overtime) did he work?

OpenStudy (anonymous):

Gimme time and Ill answer them all

OpenStudy (countryboy15):

what do you think it is

OpenStudy (anonymous):

Oh sure thing!!!

OpenStudy (countryboy15):

Since Andrew earned 330 you would substitute it for S 330 = 240 + 9y <--- solve for y 330 ( - 240 ) = 240 ( - 240 ) 9y < ---- subtract 240 from each side 90 = 9y 90 / 9 = 9y / 9 <----- divide by 9 10 =y Andrew worked 10 hours for overtime

OpenStudy (anonymous):

A) w(t) = 6 t { t< or = 40 } B) if t is greater than 40, Andrew makes 240 for is full time pay + $9 for every hour over 40 w(t) = 6 t { t< or = 40 } and = 240 + 9 (t-40) { t>40 you can simplify the second equation to w(t) = -120 + 9t {t >40 C) Andrew made 330. 330>240 so t> 40 -120 + 9t = 330 9t = 450 t = 50

OpenStudy (anonymous):

wait I messed up

OpenStudy (countryboy15):

lol

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

Take your time

OpenStudy (countryboy15):

if t is greater than 40, Andrew makes 240 for is full time pay + $9 for every hour over 40 w(t) = 6 t { t< or = 40 } and = 240 + 9 (t-40) { t>40 you can simplify the second equation to w(t) = -120 + 9t {t >40

OpenStudy (anonymous):

I did like 70 assesments today literally, I am having brainfarts!

OpenStudy (anonymous):

LOL! Thank you both very much!!!!

OpenStudy (anonymous):

nicely said CB

OpenStudy (anonymous):

Medals all around!

OpenStudy (countryboy15):

pass on medals

OpenStudy (anonymous):

Ok help me out here, how do I do that??

OpenStudy (countryboy15):

u medal need ill medal u

OpenStudy (anonymous):

Can I give you both medals?

OpenStudy (countryboy15):

nope open study only lets 1 medal per person thats why

OpenStudy (anonymous):

Ok I tried /:

OpenStudy (countryboy15):

all good did u want a medal??

OpenStudy (anonymous):

meh oly got 1 medal

OpenStudy (countryboy15):

dude it got all messed up so yeah ill give u one need